OpenStudy (anonymous):

help ,Please! A certain game has two stacks of 30 tiles with pictures on them. In the first stack of tiles, there are 10 dogs, 4 cats, 5 balls, and 11 horses. In the second stack of tiles, there are 3 flowers, 8 fish, 12 balls, 2 cats, and 5 horses. The top tile in each stack is chosen. Find each probability.

OpenStudy (anonymous):

probability of what?

OpenStudy (anonymous):

2) P (exactly one is a ball): 5/30=1/6 ?

OpenStudy (zarkon):

#2 is not correct

OpenStudy (anonymous):

exactly one is a ball. first one is + second one is - they both are

OpenStudy (anonymous):

\[\frac{1}{6}+\frac{2}{5}-\frac{1}{6}\times \frac{2}{5}\]

OpenStudy (zarkon):

P (exactly one is a ball)=P(1st is ball and 2nd is not a ball)+P(1st is not a ball and 2nd is a ball)

OpenStudy (anonymous):

i hope we get the same answer either way

OpenStudy (zarkon):

we don't

OpenStudy (anonymous):

damn then i am wrong and now i have to think

OpenStudy (zarkon):

I get \[\frac{13}{30}\]

OpenStudy (anonymous):

ok so what is the flaw in my thinking that \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] i am sure you are right, just asking

OpenStudy (zarkon):

with that formula both could be balls

OpenStudy (anonymous):

oooooooooooh yes. of course. thanks

OpenStudy (anonymous):

i guess i could subtract off twice then yes?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

of course. exclusive or. got it thanks

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

B=30-12=18 B=18/30

OpenStudy (zarkon):

for #2) doing it the way i typed above \[\frac{5}{30}\frac{18}{30}+\frac{25}{30}\frac{12}{30}\]

OpenStudy (zarkon):

you could also do this P (exactly one is a ball)=1=P(both balls or both not balls) \[=1-\frac{5}{30}\frac{12}{30}-\frac{25}{30}\frac{18}{30}\]

OpenStudy (zarkon):

1-P(both balls or both not balls)

OpenStudy (anonymous):

1) = 13/30 ?

OpenStudy (zarkon):

for #2 yes ...#1 no

OpenStudy (zarkon):

1)\[\frac{1}{15}\] 2)\[\frac{13}{30}\]